Can someone please solve this problem using system of linear equations?
Mr. White ordered pizza for the party. Medium pizzas have 8 slices and large pizzas have 10. If there are 13 pizzas and 108 slices, how many medium pizzas are there?
Any help would be appreciated.

Answer by jorel555(1290)

Let m=medium pizzas; then
8m+10(13-m)=108
8m+130-10m=108
2m=22
m=11
There are 11 medium pizzas, and 2 large.
11x8=88
+ 2x10=20==
108 total slices of pizza!!!!!!!!